# The brightest X-ray AGNs at redshift $3\lesssim z\lesssim 6$

### v1

@@ Abstract
Given recent X-ray observations of high-redshift active galactic nuclei (AGNs), we consider whether the extreme luminosities of these AGNs are consistent with current semi-analytical models. In particular, we apply extreme-value statistics (EVS) to obtain predictions of extreme X-ray luminosities of AGNs in the redshift range $3\lesssim z\lesssim 6$. We apply this formalism using different X-ray luminosity functions and compare the predicted extreme luminosities to AGNs in the Stripe-82 X-ray catalogue. We find a general consistency between data and the EVS predictions although there is some tension with certain luminosity functions. In addition, the extreme X-ray luminosities are found to be at least an order of magnitude below the Eddington luminosity limit. We discuss possible extensions to this model, including extrapolating our results to even higher redshifts ($z\gtrsim 10$) where AGNs have recently been observed.
@@
Galaxies: nuclei, statistics. X-ray: galaxies.
pubyear: 2025pagerange: The brightest X-ray AGNs at redshift $3\lesssim z\lesssim 6$A
@@ 1 Introduction
Active Galactic Nuclei (AGNs) are some of the most energetic astrophysical objects in the Universe. They are created by the accretion of matter onto supermassive black holes (SMBHs) with typical masses of $M_{\rm BH}\sim 10^{5}$$10^{9}$ $M_{\odot}$ at their cores (Rees, 1984; Fabian, 2012; Yuan & Narayan, 2014; Netzer, 2015; Inayoshi et al., 2020). Among the various wavelengths used to study AGNs, X-ray observations provide a unique window into the innermost regions closest to the event horizon of the black hole.
@@
The most luminous AGNs, with X-ray luminosity $L_{\rm X}$ as large as $10^{45}$ erg s-1, stand out as extraordinary laboratories for understanding extreme astrophysical processes (e.g. Stanley et al., 2017; Veronesi et al., 2023). Their immense luminosities result from the highly efficient energy conversion in the accretion process, making them critical probes of the interplay between SMBHs and their extreme environments.
@@
Our EVS modelling will allow us to test various semi-analytical models of the X-ray luminosity functions by comparing their predictions of extreme X-ray luminosities against data from the Stripe 82 X-ray catalogue, a benchmark dataset for X-ray surveys combining data from the Chandra and XMM-Newton space telescopes (LaMassa et al., 2013, 2016, 2024). In particular, we will use the so-called S82-XL data from Peca et al. (2024) and LaMassa et al. (2024), and focus on AGNs at redshifts $3\lesssim z\lesssim 6$. The central goal of this work is to quantify the consistency (or tension) between theory and observation.
The rest of this paper is organized as follows. In Section 2, we present a survey of models of the X-ray luminosity functions and their associated observables. Section 4 details our EVS modelling, with the main results (the predictions of of extreme AGN luminosities across redshifts) shown in Section 5. We discuss the implications of our findings in Section 6 .
We will assume a flat $\Lambda\rm{CDM}$ cosmology with $H_{0}=70\,\rm{kms^{-1}Mpc^{-1}}$, $\Omega_{m}=0.3$ and $\Omega_{\Lambda}=0.7$. We work in natural units with $c=1$.
@@ 2 The X-Ray Luminosity Function
Figure 1: Comparison of the X-ray luminosity functions at $z=3$ (left), $z=4.5$ (middle) and $z=6$ (right). We include the PDE and LDDE model from Pouliasis et al. (2024), the LDDE models from Georgakakis et al. (2015) and Vito et al. (2014), as well as a LDDE with a redshift cut-off from Ueda et al. (2014).
@@
In this section, we explore two forms of $e(z)$, namely the pure density evolution (PDE) model and the luminosity-dependent density evolution (LDDE) model. We also examine a form of the LDDE model with a cut-off redshift value. Fig. 1 gives a graphical summary of the various luminosity functions explored in this section.
@@ 2.1 Pure density evolution (PDE) model
Fitting this to AGN data at redshifts 33336666, Pouliasis et al. (2024) gave the value $p_{\rm den}=-7.35^{+0.39}_{0.42}$ with $z_{c}=3$. Assuming the PDE model, the parameter values in the luminosity function equation (1) are
@@ 2.2 Luminosity-dependent density evolution (LDDE) model
We will consider various parametrisations of this form given in Pouliasis et al. (2024), Georgakakis et al. (2015) and Vito et al. (2014). The associated parameter values from these references are shown in Table 1.
@@ 2.3 LDDE model with redshift cut-off
This is a model with a more complicated evolution factor of the form (Ueda et al., 2014)
$e(z,L_{\rm X})=\begin{cases}(1+z)^{p_{1}}&[z\leq z_{c_{1}}(L_{\rm X})]\\ (1+z_{c_{1}})^{p_{1}}\left(\frac{1+z}{1+z_{c_{1}}}\right)^{p_{2}}&[z_{c_{1}}(L_{\rm X})<z\leq z_{c_{2}}]\\ (1+z_{c_{1}})^{p_{1}}\left(\frac{1+z}{1+z_{c_{1}}}\right)^{p_{2}}\left(\frac{1+z}{1+z_{c_{2}}}\right)^{p_{3}}&[z>z_{c_{2}}],\end{cases}$ (5)
@@
$z_{c_{1}}(L_{X})=\begin{cases}z^{*}_{c1}(L_{\rm X}/L_{a_{1}})^{\alpha_{1}}&[L_{\rm X}\leq L_{a_{1}}]\\ z^{*}_{c_{1}}&[L_{\rm X}>L_{a_{1}}],\end{cases}$ (7)
@@
$z_{c_{2}}(L_{X})=\begin{cases}z^{*}_{c_{2}}(L_{\rm X}/L_{a_{2}})^{\alpha_{2}}&[L_{\rm X}\leq L_{a_{2}}]\\ z^{*}_{c_{2}}&[L_{\rm X}>L_{a_{2}}].\end{cases}$ (8)
@@ 3 AGN number count
For an AGN with X-ray luminosity $L_{X}$ at redshift $z$, we also consider its hydrogen column density $N_{\rm H}$, which quantifies how obscured the AGN is. Higher $N_{\rm H}$ values indicate significant obscuration, for example, due to absorption by a dusty torus or host-galaxy gases (Antonucci, 1993; Netzer, 2015). Compton-thick AGNs with $N_{\rm H}>10^{24}{\rm cm}^{-2}$ can completely obscure soft X-ray emissions. Understanding $N_{\rm H}$ is essential for probing the physical properties of AGN environments and their impact on observed spectra (Elitzur & Shlosman, 2006).
@@
×$\times f_{\rm sky}\frac{\,\textrm{d}V}{\,\textrm{d}z}\,\textrm{d}\log L_{\rm X}\ \,\textrm{d}z\ \,\textrm{d}\log N_{\rm H},$
@@
$H(z)=H_{0}\left[\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}\right]^{\frac{1}{2}}.$ (11)
The absorption function $f_{\rm abs}$ depends on X-ray luminosity, redshift and $N_{\rm H}$. We use the functional form of $f_{\rm abs}$ given in Ueda et al. (2003), and choose $f_{\rm sky}=54.8\ {\rm deg}^{2}$, in agreement with the S82-XL survey.
@@ 4.1 Extreme-value statistics
We utilise the generalised extreme-value (GEV) method, also known as the block maxima method (Gumbel, 1958; de Haan & Ferreira, 2006). In this method, we divide the AGN population in a given redshift bin into $N$ distinct blocks, from each of which we identify the brightest AGN. Analogous to the Central Limit Theorem, in the large-$N$ limit, the distribution of the brightest AGNs will approach the generalised extreme-value distribution (equation (14)).
@@
$\gamma&=N(>L_{\rm peak})-1,\\ \\ \alpha&=L_{\rm peak}-\frac{\beta}{\gamma}\left({(1+\gamma)}^{-\gamma}-1\right),\end{aligned}\begin{aligned} \beta&=\frac{{(1+\gamma)}^{1+\gamma}}{\displaystyle\frac{\,\textrm{d}N}{\,\textrm{d}L_{\rm X}}\Big{|}_{L_{\rm peak}}}.\end{aligned}$ (15)
In Fig. 2, we plot the extreme-value pdf for the brightest AGNs for three redshift bins $z=[3,3.5],[4.5,5]$ and $[5,5.5]$ where the distributions shown are calculated from equation (14). The values for $\alpha$ and $\beta$, are calculated from equations 15 and shown in Table 2 in the Appendix. We note from the Table that $\alpha$ is indeed close to the peak luminosity in each redshift bin.
@@
Figure 3: A plot to show the comparison of the Gumbel fit to the Poisson pdf with three redshift bins for the cut-off redshift LDDE luminosity function in Section 2.3. The solid coloured lines show the pdf derived from the Poisson distribution (equation 13) and the dashed lines are the Gumbel distribution (equation 14) with $\gamma=0$ and the values for $\alpha$ and $\beta$ given in Table 2. We can see that choosing the Gumbel distribution, where we take $\gamma=0$, is a suitable fit for our data.
@@ 4.2 Stripe 82 X-ray AGN data
The Stripe 82 X-ray survey is a comprehensive astronomical survey focused on detecting rare, high-luminosity AGNs and other X-ray sources within the area covered by the Stripe 82 field of the Sloane Digital Sky Survey (SDSS). The survey incorporates data from the XMM-Newton and Chandra telescopes to create a comprehensive catalogue of X-ray sources (e.g. LaMassa et al., 2013, 2016). The continuous updates to the catalogue have been invaluable to the study of the high-redshift Universe.
We take data from the S82X catalogues as described in LaMassa et al. (2024) and Peca et al. (2024). These catalogues provide samples at $z\lesssim 3.5$ and $3\lesssim z\lesssim 6$, with sky survey areas of $\sim 31.3$ deg2 and 54.8 deg2, respectively.
LaMassa et al. (2024) also gave the Eddington luminosity ($L_{\rm Edd}$) for samples with $z\lesssim 3.5$, which was calculated using the reference black hole mass according to the formula
$L_{\rm Edd}=1.26\times 10^{38}~{}M_{\rm BH}/M_{\odot}~{}\rm{erg}~{}\rm{s}^{-1}.$ (16)
[…]
For the EVS analysis, we will use a subset of 35 AGNs from the S82X data. This is comprised of 17 sources from LaMassa et al. (2024) in the redshift range $3\leq<z<\leq$ which where given values for the Eddington limit, plus the brightest 18 sources from Peca et al. (2024) in the range $3.5\leq z\leq 6$.
Figure 4: The peak luminosities of $L_{\rm X}$ (solid black lines) of the extreme-value (Gumbel) pdf for the $L_{\rm X}$ functions from Section 2, plotted in redshift bins of size $0.5$ from $z=3$ to $z=6$. Each block shows a profile of the EVS probability density function shown in fig. 2 where the darker/ lighter shaded regions correspond to the $95^{th}$ and $99^{th}$ percentiles. The data points (in green) are from the S82-XL survey Peca et al. (2024), and the Eddington luminosities (blue) are from LaMassa et al. (2024).

### v2

@@ Abstract
Given recent X-ray observations of high-redshift active galactic nuclei (AGNs), we consider whether the extreme luminosities of these AGNs are consistent with current semi-analytical models. In particular, we apply extreme-value statistics (EVS) to obtain predictions of extreme X-ray luminosities of AGNs in the redshift range $3\lesssim z\lesssim 6$. We apply this formalism using different X-ray luminosity functions and compare the predicted extreme luminosities to AGNs in the Stripe 82 X-ray catalogue. We find a general consistency between data and the EVS predictions although there is some tension with certain luminosity functions. We discuss possible extensions to this model, including extrapolating our results to even higher redshifts ($z\gtrsim 10$) where AGNs have recently been observed.
@@
Galaxies: nuclei, statistics. X-ray: galaxies. Methods: statistical.
pubyear: 2025pagerange: The brightest X-ray AGNs at redshift $3\lesssim z\lesssim 6$B
@@ 1 Introduction
Active Galactic Nuclei (AGNs) are some of the most energetic astrophysical objects in the Universe. They are created by the accretion of matter onto supermassive black holes (SMBHs) with typical masses of $M_{\rm BH}\sim 10^{5}$$10^{9}$ $\rm M$$_{\odot}$ at their cores (Rees, 1984; Fabian, 2012; Yuan & Narayan, 2014; Netzer, 2015; Inayoshi et al., 2020). Among the various wavelengths used to study AGNs, X-ray observations provide a unique window into the innermost regions closest to the event horizon of the black hole.
@@
The most luminous AGNs, with 2–10 keV X-ray luminosity $L_{\rm X}$ as large as $10^{45}$ erg s-1, stand out as extraordinary laboratories for understanding extreme astrophysical processes (e.g. Mateos et al., 2015; Stanley et al., 2017; Veronesi et al., 2023). Their immense luminosities result from the highly efficient energy conversion in the accretion process, making them critical probes of the interplay between SMBHs and their extreme environments.
@@
Our EVS modelling will allow us to test various semi-analytical models of the X-ray luminosity functions by comparing their predictions of extreme X-ray luminosities against data from the Stripe 82 X-ray catalogue, a benchmark dataset for X-ray surveys combining data from the Chandra and XMM-Newton space telescopes (LaMassa et al., 2013, 2016, 2024). In particular, we will use the so-called S82-XL data from Peca et al. (2024) and LaMassa et al. (2024), and focus on AGNs at redshifts $3\lesssim z\lesssim 6$. The central goal of this work is to quantify the consistency (or tension) between theory and observation. The EVS formalism does not require a complete census of all AGNs between $3<z<6$ in Stripe 82X. Instead, we only require the X-ray luminosities of the most luminous AGNs observed in LaMassa et al. (2024) and Peca et al. (2024). In other words, our focus is on comparing the brightest, confirmed AGNs observed in Stripe 82X to the upper bound of AGN luminosities predicted by EVS.
The rest of this paper is organized as follows. In Section 2, we present a survey of models of the X-ray luminosity functions and their associated observables. In Section 3, we present the formalism used to calculate AGN number counts. Section 4 details our EVS modelling, with the main results (the predictions of of extreme AGN luminosities across redshifts) shown in Section 5. We discuss the implications of our findings in Section 6 .
We will assume a flat $\Lambda\rm{CDM}$ cosmology with $H_{0}=70\,\rm{km\ s^{-1}\ Mpc^{-1}}$, $\Omega_{\rm m}=0.3$ and $\Omega_{\Lambda}=0.7$. We work in natural units with $c=1$.
@@ 2 The X-Ray Luminosity Function
Figure 1: Comparison of the X-ray luminosity functions at $z=3$ (left), $z=4.5$ (middle) and $z=6$ (right). We include the PDE and LDDE model from Pouliasis et al. (2024), the LDDE models from Georgakakis et al. (2015) and Vito et al. (2014), a LDDE with a redshift cut-off from Ueda et al. (2014), and the Peca et al. (2023) model derived from numerical fit to observation.
@@
In this section, we explore two forms of $e(z)$, namely the pure density evolution (PDE) model and the luminosity-dependent density evolution (LDDE) model. The two models differ in their dependence on X-ray luminosity. In the PDE model, we assume that the redshift evolution of the number density is the same regardless of AGN luminosity. In contrast, the LDDE model assumes that the redshift evolution depends on the luminosity of the AGNs. Although the LDDE model includes an additional parameter, it provides a better fit to the data, and its extrapolation below the survey flux limit remains consistent with the cosmic X-ray background (Miyaji et al., 2000). We also examine a form of the LDDE model with a cut-off redshift value. Fig. 1 gives a graphical summary of the various luminosity functions explored in this section.
@@ 2.1 Pure density evolution (PDE) model
The parameter $p_{\rm den}$ gives the slope of the power law dependence on redshift $z$. $z_{c}$ is the critical redshift. Fitting this to AGN data at redshifts 3366, Pouliasis et al. (2024) gave the value $p_{\rm den}=-7.35^{+0.39}_{-0.42}$ with $z_{c}=3$. Assuming the PDE model, the parameter values in the luminosity function equation (1) are
@@ 2.2 Luminosity-dependent density evolution (LDDE) model
$\beta$ is the slope of the luminosity-dependence redshift evolution while $p_{\rm den}$ is the slope of the power law redshift evolution similar to equation (3). We will consider various parametrisations of this form given in Pouliasis et al. (2024), Georgakakis et al. (2015) and Vito et al. (2014). The associated parameter values from these references are shown in Table 1.
@@ 2.3 LDDE model with redshift cut-off
This is a model with a more complicated evolution factor of the form (Ueda et al., 2014). This model introduces additional redshift cut-offs, $z_{c_{1}}$ and $z_{c_{2}}$, as well as slopes of the power law; $p_{1},p_{2}$ and $p_{3}$:
$e(z,L_{\rm X})=\begin{cases}(1+z)^{p_{1}}&[z\leq z_{c_{1}}(L_{\rm X})]\\[8.61108pt] (1+z_{c_{1}})^{p_{1}}\left(\displaystyle\frac{1+z}{1+z_{c_{1}}}\right)^{p_{2}}&[z_{c_{1}}(L_{\rm X})<z\leq z_{c_{2}}]\\[8.61108pt] (1+z_{c_{1}})^{p_{1}}\left(\displaystyle\frac{1+z}{1+z_{c_{1}}}\right)^{p_{2}}\left(\displaystyle\frac{1+z}{1+z_{c_{2}}}\right)^{p_{3}}&[z>z_{c_{2}}],\end{cases}$ (5)
@@
$z_{c_{1}}(L_{\rm X})=\begin{cases}z^{*}_{c_{1}}(L_{\rm X}/L_{a_{1}})^{\alpha_{1}}&[L_{\rm X}\leq L_{a_{1}}]\\[8.61108pt] z^{*}_{c_{1}}&[L_{\rm X}>L_{a_{1}}],\end{cases}$ (7)
@@
$z_{c_{2}}(L_{\rm X})=\begin{cases}z^{*}_{c_{2}}(L_{\rm X}/L_{a_{2}})^{\alpha_{2}}&[L_{\rm X}\leq L_{a_{2}}]\\[8.61108pt] z^{*}_{c_{2}}&[L_{\rm X}>L_{a_{2}}].\end{cases}$ (8)
@@
2.4 Numerical fit model
We have also included the luminosity function from Peca et al. (2023), which is a numerical fit to the observed $L_{\rm X}$ values from the Stripe 82X catalogue up to $z=4$. We numerically extrapolated the results to $z=6$ using a linear grid interpolator on the $L_{\rm X}$-$z$ plane; the data is fitted as a surface on this rectilinear grid and values outside this domain can be extrapolated, just as the authors have done in their work.
@@ 3 AGN number count
For an AGN with X-ray luminosity $L_{\rm X}$ at redshift $z$, we also consider its hydrogen column density $N_{\rm H}$, which quantifies how obscured the AGN is. Higher $N_{\rm H}$ values indicate significant obscuration, for example, due to absorption by a dusty torus or host-galaxy gases (Antonucci, 1993; Netzer, 2015). Compton-thick AGNs with $N_{\rm H}>10^{24}\,{\rm cm}^{-2}$ can completely obscure soft X-ray emissions. Understanding $N_{\rm H}$ is essential for probing the physical properties of AGN environments and their impact on observed spectra (Elitzur & Shlosman, 2006).
@@
$\times f_{\rm sky}\frac{\,\textrm{d}V}{\,\textrm{d}z}\,\textrm{d}\log L_{\rm X}\ \,\textrm{d}z\ \,\textrm{d}\log N_{\rm H},$
@@
$H(z)=H_{0}\left[\Omega_{\rm m}(1+z)^{3}+\Omega_{\Lambda}\right]^{\frac{1}{2}}.$ (11)
The absorption function $f_{\rm abs}$ depends on X-ray luminosity, redshift and $N_{\rm H}$. While Ueda et al. (2014) gave an updated XLF and absorption function built into a population synthesis model, we opted to use the functional form of $f_{\rm abs}$ from Ueda et al. (2003) because it provides a simple analytic prescription that is widely used in XLF modelling, enabling a more direct comparison with prior works. The differences between both models are minimal at the high X-ray luminosities that dominate our samples. The Compton-thick AGNs ($N_{\rm H}>10^{24}~{\rm cm}^{-2}$) which are not explicitly included in our $f_{\rm abs}$ integration, but are more carefully treated in Ueda et al. (2014), are not included in our sample. In particular, there are no Compton-thick AGN reported in the catalogue in the redshift range $3\leq z\leq 6$. Indeed, the observed Stripe 82X catalogue showed a median $N_{\rm H}=10^{21.6}~{\rm cm}^{-2}$, with only  37% of sources having $N_{\rm H}>10^{22}~{\rm cm}^{-2}$ (Peca et al., 2024). In other words, most of the brightest AGNs in our sample are essentially unobscured, so omitting Compton-thick AGNs is expected to have only a negligible effect on our results.
The number count in equation (9) requires intrinsic X-ray luminosity, whereas the AGN from our sample are given in observed X-ray luminosity. We find the difference between the observed and intrinsic luminosity to be negligible, as verified using PIMMS [^1] https://heasarc.gsfc.nasa.gov/docs/software/tools/pimms.html under the standard spectral assumptions for unobscured AGNs (e.g. $\Gamma=1.8$). We found that even at the higher column densities in our sample ($N_{\rm H}\approx 10^{23.5}\rm\ cm^{-2}$), the percentage difference between the intrinsic and observed luminosities is still relatively small ($\sim 10\%$ for $z=3$, decreasing to $\sim 2.5\%$ for $z=6$). The percentage difference could be higher for Compton-thick AGNs but these are not present in our data.
The S82-XL survey reaches very faint X-ray fluxes (hard-band 2–10 keV limit $\sim 2.9\times 10^{-15}$ erg s-1 cm-2). We found that 34 of our 35 sources are above the flux limit, with almost all having a flux greater than $10^{14}\rm erg\ s^{-1}\ cm^{-2}$. There is only one source we are using that is below this flux limit, which corresponds to the AGN at $z=5.855$, previously studied elsewhere Pâris et al. (2012). We chose to keep this point as it has the highest spectroscopically confirmed redshift in the catalogue, which is important in our EVS discussion.
Although the sky coverage can vary with flux sensitivity (introducing a bias against Compton-thick AGNs, which appear fainter), this effect is negligible for our sample of bright AGNs, the majority of which are also Compton-thin and detectable across the entire survey area. We therefore adopt a constant $f_{\rm sky}=54.8$ deg2 in our analysis, as in Peca et al. (2024). Other factors, such as the incompleteness of spectroscopic redshifts, may also directly affect the expected maximum luminosity from the XLF models. We discuss an uncertainty factor, $f_{\rm unc}$, which modulate the number counts, in Appendix A. In any case, we find that this has a subdominant effect on our extreme-value distributions and does not change our main conclusions.
@@ 4.1 Extreme-value statistics
We utilise the generalised extreme-value (GEV) method, also known as the block maxima method (Gumbel, 1958; de Haan & Ferreira, 2006). In this method, we divide the AGN population in a given redshift bin into $S$ distinct blocks, from each of which we identify the brightest AGN. Analogous to the Central Limit Theorem, in the large-$S$ limit, the distribution of the brightest AGNs will approach the generalised extreme-value distribution (equation (14)).
@@
$\begin{aligned} \gamma&=N(>L_{\rm peak})-1,\\ \\ \alpha&=L_{\rm peak}-\frac{\beta}{\gamma}\left({(1+\gamma)}^{-\gamma}-1\right),\end{aligned}\begin{aligned} \beta&=\frac{{(1+\gamma)}^{1+\gamma}}{\displaystyle\frac{\,\textrm{d}N}{\,\textrm{d}L_{\rm X}}\Big|_{L_{\rm peak}}}.\end{aligned}$ (15)
tIn Fig. 2, we plot the extreme-value pdf for the brightest AGNs for three redshift bins $z=[3,3.5],[4.5,5]$ and $[5,5.5]$ where the distributions shown are calculated from equation (14). The values for $\alpha$ and $\beta$, are calculated from equations (15) and shown in Table 2 in the Appendix B. We note from Table 2 that $\alpha$ is indeed close to the peak luminosity in each redshift bin.
@@
Figure 3: A plot to show the comparison of the Gumbel fit to the Poisson pdf with three redshift bins for the cut-off redshift LDDE luminosity function in Section 2.3. The solid coloured lines show the pdf derived from the Poisson distribution (equation (13)) and the dashed lines are the Gumbel distribution (equation (14)) with $\gamma=0$ and the values for $\alpha$ and $\beta$ given in Table 2. We can see that choosing the Gumbel distribution, where we take $\gamma=0$, is a suitable fit for our data.
@@ 4.2 Stripe 82 X-ray AGN data
The Stripe 82 X-ray survey is a comprehensive astronomical survey focused on detecting rare, high-luminosity AGNs and other X-ray sources within the area covered by the Stripe 82 field of the Sloan Digital Sky Survey (SDSS). The survey incorporates data from the XMM-Newton and Chandra telescopes to create a comprehensive catalogue of X-ray sources (e.g. LaMassa et al., 2013, 2016). The continuous updates to the catalogue have been invaluable to the study of the high-redshift Universe.
We take data from the S82X catalogue as described in Peca et al. (2024). This catalogues provides samples of AGN at redshift values $3\lesssim z\lesssim 6$, with a sky survey area of 54.8 deg2. We are using the 2–10 keV observed luminosity from the catalogue as a proxy for $L_{\rm X}$.
For the EVS analysis, we use a subset of 35 most luminous AGNs from the S82X data from Peca et al. (2024) that were in the redshift range $3\lesssim z\lesssim 6$, and were above a certain threshold luminosity. In the catalogue there is a large number of sources with redshift $3\lesssim z\lesssim 3.5$ with luminosity $44\lesssim\log(L_{\rm X})\lesssim 45$. We chose to restrict the number of sources to only the brightest ones in this redshift range.
Figure 4: The peak luminosities of $L_{\rm X}$ (solid black lines) of the extreme-value (Gumbel) pdf for the $L_{\rm X}$ functions from Section 2, plotted in redshift bins of size $0.5$ from $z=3$ to $z=6$. Each block shows a profile of the EVS probability density function shown in Fig. 2 where the darker/ lighter shaded regions correspond to the $95^{th}$ and $99^{th}$ percentiles. The data points (green) present the most luminous AGN in the S82-XL survey from Peca et al. (2024).